the sum of two numbers is 91.
12 times the smaller number is 32 more than 8 times the larger.
find the numbers

Answer by mananth(16949)

the sum of two numbers is 91.
12 times the smaller number is 32 more than 8 times the larger.
find the numbers

the sum of two numbers is 91. let x < y
x+y=91............................(1)
12 times the smaller number is 32 more than 8 times the larger.
12x=32+8y
12x-8y=32
/4
3x-2y=8............................(2)

1 x + 1 y = 91 .............1

3 x -2 y = 8 .............2
Eliminate y
multiply (1)by 2
Multiply (2) by 1
2 x 2 y = 182
3 x -2 y = 8
Add the two equations
5 x = 190
/ 5
x = 38
plug value of x in (1)
1 x + 1 y = 91
38 + y = 91
y = 91 -38
y = 53
y = 53 1
x= 38
y= 53
m.ananth@hotmail.ca